All posts by Jeremiah Tsyporin

ISB Observations 8/22

Jeremiah went to the ISB roof this morning. A 75mL sample (S-UCSC-140822) was present, and a blank (B-UCSC-140822) was taken after cleaning the collector. The collector is in very good condition. The sample collected this morning contained a thin layer of what looked like dirt or soil at the bottom.


Jeremiah went to the roof of ISB this morning. The fog collector was not running. RH sensor is currently not working and the collector needs to be manually turned on the night before collection. Most of the days this week are confirmed for somebody turning on the collector the night before however communication errors are responsible for the lack of sample this morning. I will be turning on the collector Wednesday night.

140626 UCSC Observations


This morning the fog collector on the roof of ISB was open and running. There was no sample. I dismantled and cleaned the fog collector then took a blank (B-UCSC-140626). The collector is in very good condition, some spiders were found inside however it was overall clean. Log book has been returned to core lab and is in the bucket with the keys to ISB.